OpenStudy (anonymous):

What is the factorization of the trinomial below? 3x2 + 27x + 60

OpenStudy (anonymous):

factor a 3 what do you get?

OpenStudy (anonymous):

3(x^2+9x+20)

OpenStudy (anonymous):

ok now what are the 2 numbers that there product is 20 and sum is 9

OpenStudy (anonymous):

3 and 2 ?

OpenStudy (anonymous):

3*2=?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

you need 20 so what are the factors of 20?

OpenStudy (anonymous):

@ermete.mayas u there

OpenStudy (anonymous):

They are 1,2,4,5,10.

OpenStudy (anonymous):

which 2 will sum up to 9?

OpenStudy (anonymous):

lol im lost

OpenStudy (anonymous):

4*5=? 4+5=?

OpenStudy (anonymous):

20, 9

OpenStudy (anonymous):

correct 3(x^2+9x+20) now 3(x+5)(x+4)

OpenStudy (anonymous):

thank you